OpenStudy (anonymous):

I was given a circle with the radius of x+2 and the circumference of 31.40. Find the value of x.

OpenStudy (sandra):

right, so the formula for the circumference of a circle is: \[c = \pi*d\] where d is the diameter of the circle (aka 2 times the radius)

OpenStudy (sandra):

so we have: \[31.40 = \pi*(x+2/2)\] \[31.40 = (\pi*x + 2*\pi)/2\]

OpenStudy (sandra):

actually sorry, easier to say : 31.40 = (pi * (x+2))/2 62.80 = pi*(x+2) x+2 = 62.80/pi And you have to finish the rest =)